# Request-Only Optimization for Recommendation Systems

### v1

@@ 5. Related Work
A previous design RecD (Zhao et al., 2023) addressed the feature duplication problem by implementing a suite of data pipeline optimizations to group impression samples of the same session so that columnar storage compression ratio could be maximized. It introduced an inverse index to reconstruct the impression training samples for feature preprocessing, sparse data distribution, and embedding lookups during model training. Our work differs from RecD with a a holistic approach to ensure systematic efficiency across storage, training, and inference. ROO data schema deduplicates feature at the root source of the training data pipeline, as a result it does not entail data pipeline infrastructure changes nor does it increase data landing latency. Its feature preprocessing integrates seamlessly with commonly used tensor data structures (e.g. KeyedJaggedTensors in TorchRec (Ivchenko et al., 2022)) without any custom data structures like InverseKeyedJaggedTensors, minimizing potential subsequent model rewrite. In appendix, we explain ROO data with more implementation details.
Efficient scaling of DLRM. To scale DLRMs, an effective line of thinking is to examine the recommendation life-cycle from an end-to-end point of view, seeking model-system co-design and global optimization opportunities. Efficient application of various hardware (e.g. accelerators, scheduler, SSDs) in different components of the end-to-end recommendation system has been studied in (Mudigere et al., 2022; Gupta et al., 2020a; Wilkening et al., 2021). In particular, some work adapted the ML strategy to enable more efficient usages of the accelerator resources (Xu et al., 2022; Zhai et al., 2023). In addition, leveraging offline/batch compute that has access to more compute and memory resources is also observed in many systems (Pancha et al., 2022; Pi et al., 2019).
@@
Being orthogonal to the innovative methods above, the ROO paradigm offers a generic solution from the angle of systematic de-duplication and amortization. This approach underscores the importance of model and system co-design in jointly optimizing data, training and inference processes under the DLRM context.
@@ 7. Acknowledgment
This work would not be possible without work from the following contributors (alphabetical order): Bugra Akyildiz, Yuanfei Bi, Duo Chen, Li Chen, Xi Chen, Xianjie Chen, Rex Cheung, Patrick Cullen, Delia David, Fei Ding, Rajeev Dixit, Lu Fang, Md Mustafijur Rahman Faysal, Bo Feng, Liz Guo, Luyi Guo, Daisy Shi He, Michael He, Austin Jenkins, Jie Jiao, Basri Kahveci, Manos Karpathiotakis, Keyu Lai, Dai Li, Shen Li, Hao Lin, Weiran Liu, Yinghai Lu, Michael Ly, Matt Ma, Yun Mao, Franco Mo, Min Ni, Jongsoo Park, Jing Qian, Alex Singh, Jason Song, Bingjun Sun, Ruohan Sun, Yimin Tan, Alex Vu, Shanyue Wang, Wenyu Wang, Zellux Wang, Ziqi Wang, Evan Welch, Yue Weng, Yanhong Wu, Ke Xu, Zheng Yan, Zimeng Yang, James Zhang, Jingjing Zhang, Lu Zhang, Qunshu Zhang, Vivian Zhang, Wei Zhang, Yue Zhang, Zhiyun Zhang, Xin Zhuang.
@@ Appendix A Notations
$num\_candidate$ number of candidates in a request
@@
$MT$ Multi-task prediction for ranking models.
@@ Appendix B Request Level Join
  • Due to feature deduplication at the root data source of the training data pipeline, the ROO training data does not require additional data infra feature, nor does it add any data landing latency increase. In contrast, RecD requires extra ETL jobs to cluster samples by session, which adds data landing latency by 30 minutes in production and makes it inherently incompatible for online training and model responsiveness.
@@
In one recommendation product surface, we replaced a RecD-trained LSR model with ROO-trained HSTU model and ROO training showed a 10% training QPS win even without any overarch optimization. On the same product surface, the ROO trained ESR model showed 50% training QPS win over the RecD version.
@@
4:Set¡Int¿ impressions;$\triangleright$ impression items
5:Map¡Int, List¡Int¿¿ conversions;$\triangleright$ item labels
@@
16:     if $userHasNewRequestId(requestId,currRequestId)$ —— $exceedsUserEngagementThreashold(joinKey)$ —— $joinWindowTimeUp(joinKey)$ then
17:         closeJoinWindow(joinKey)$\triangleright$ publish ROO sample
@@
21:     if $shouldStartNewJoinWindow$ then
22:         rec = RequestJoinRecord($\triangleright$ new join record
@@
25:         Joiner.set(JoinKey($userId,requestId$), $rec$)
@@
27:         rec = Joiner.get($joinKey$$\triangleright$ update join record
@@
Appendix C Ranking Model Architecture
@@
The late stage ranking model architecture. HSTU and user arch (highlighted) introduced RO computations.

### v3

@@ 5. Related Work
RecD (Zhao et al., 2023) addressed feature duplication through data pipeline optimizations, grouping impression samples to maximize columnar storage compression and using an inverse index for reconstruction during training. Our approach differs by offering a holistic system-wide efficiency across storage, training, and inference. The ROO data schema deduplicates features at the source, avoiding data pipeline infrastructure changes or increased landing latency. Its preprocessing integrates seamlessly with common tensor structures (e.g., KeyedJaggedTensors in TorchRec (Ivchenko et al., 2022)), eliminating the need for custom structures like InverseKeyedJaggedTensors and minimizing model rewrite efforts. More implementation details on ROO data are provided in the appendix B.
Efficient scaling of DLRM. To scale DLRMs, an effective line of thinking is to examine the recommendation life-cycle from an end-to-end point of view, seeking model-system co-design and global optimization opportunities. Efficient application of various hardware (e.g. accelerators, scheduler, SSDs) in different components of the end-to-end recommendation system has been studied in (Mudigere et al., 2022; Gupta et al., 2020a; Wilkening et al., 2021). In particular, some work adapted the ML architectures deliberately to enable more efficient usages of the accelerator resources (Xu et al., 2022; Zhai et al., 2023). In addition, leveraging offline/batch compute that has access to more compute and memory resources is also observed in many systems (Pancha et al., 2022; Pi et al., 2019).
@@
Orthogonal to the methods above, the ROO paradigm provides a generic solution focused on systematic de-duplication and amortization, highlighting the importance of model-system co-design to optimize data, training, and inference in the DLRM context.
@@ 7. Acknowledgment
This work would not be possible without work from the following contributors (alphabetical order): Bugra Akyildiz, Yuanfei Bi, Duo Chen, Li Chen, Xi Chen, Xianjie Chen, Rex Cheung, Ted Cui, Patrick Cullen, Delia David, Fei Ding, Rajeev Dixit, Md Mustafijur Rahman Faysal, Bo Feng, Liz Guo, Luyi Guo, Daisy Shi He, Michael He, Austin Jenkins, Jie Jiao, Basri Kahveci, Manos Karpathiotakis, Keyu Lai, Dai Li, Shen Li, Hao Lin, Weiran Liu, Yinghai Lu, Michael Ly, Matt Ma, Yun Mao, Franco Mo, Min Ni, Jongsoo Park, Jing Qian, Alex Singh, Jason Song, Bingjun Sun, Ruohan Sun, Yimin Tan, Alex Vu, Shanyue Wang, Wenyu Wang, Zellux Wang, Ziqi Wang, Evan Welch, Yue Weng, Yanhong Wu, Ke Xu, Zheng Yan, Zimeng Yang, James Zhang, Jingjing Zhang, Lu Zhang, Qunshu Zhang, Vivian Zhang, Wei Zhang, Yue Zhang, Zhiyun Zhang, Xin Zhuang.
@@ Appendix A Notations
num_candidatenum\_candidate number of candidates in a request
@@
MTMT Multi-task prediction for ranking models.
@@ Appendix B Request Level Join
  • Due to feature deduplication at the root data source of the training data pipeline, the ROO training data does not require additional data infra feature, nor does it add any data landing latency increase. In contrast, RecD requires extra ETL jobs to cluster samples by session, which adds data landing latency by 30 minutes in production and makes it inherently incompatible for online training and model responsiveness. ROO data has a notable advantage over RecD.
@@
[…]
@@
16:  if userHasNewRequestId(requestId,currRequestId)userHasNewRequestId(requestId,currRequestId) —— $exceedsUserEngagementThreashold(joinKey)$ —— $joinWindowTimeUp(joinKey)$ then
[…]
@@
21:  if shouldStartNewJoinWindowshouldStartNewJoinWindow then
[…]
@@
25:   Joiner.set(JoinKey($userId,requestId$), recrec)
@@
[…]
@@
Appendix C ROO Expansion for Backward Compatibility
Despite its significant storage efficiency advantage, request-level training sample schema is incompatible with the traditional DLRMs trained with impression-level schema. In a typical industrial environment, hundreds of models across multiple ranking stages collectively serve billions of users using a product, and not all models have to adopt the ROO training paradigm due to return-on-investment (ROI) considerations. We present a practical way to minimize the migration cost for these models and facilitate the adoption of ROO data schema. We introduce ROO expansion data adapter in the feature preprocessing layer.
For each mini-batch, the data adapter expands a ROO sample into multiple in-memory impression samples, with user-side features duplicated to the in-memory data frames. Trainer side rebatching is leveraged to keep the consistent training batch size. This equips ROO data schema with backward compatibility and eliminates unnecessary model rewrite efforts. Note that the ROO expansion requires more memory and computation cost on the data preprocessing workers. However, reading ROO samples has a 10x drop IO bytes from the warm storage (Hive) as compared to reading impression samples. Despite the expansion overhead, it had a 2-3x data preprocessing throughput increase. The expansion functionality simplifies the migration process; major products fully migrated all retrieval, early stage ranking and late stage ranking models to ROO data within a 9-month time frame.
Appendix D Ranking Model Architecture
@@
The late stage ranking model architecture. HSTU and user arch (highlighted) introduced additional RO computations.